Grade 12General Physics

A particle moves along a circular path of radius 20cm with a constant angular acceleration of 4 rad per sec then the angular displacement of the particle after 5 sec will be

Himanshu Vaidya
8 Years agoGrade 12

1 Answer

Aditya Mitra
8 Years ago
Answer:1.16 rad
 
He moves 5*4=20rad
 
Distance moved=angle*radius=20*20=400cm
 
Number of complete turns= dist/(2*pi*r)
=400/(2*3.14*20)=3 complete turns
 
Left over distance =total distance - distance in 3 rounds
=400-(3*2*3.14*20)
=400-376.8
=23.2 cm
 
Angular displacement =s/r
=23.2/20
=1.16 rad